OpenStudy (anonymous):

After 13 hours of burning, a candle has a height of 25.4 centimeters. After 24 hours of burning its height is 23.2 centimeters. What is the height after 17 hours of burning?

OpenStudy (anonymous):

Let x be the original height \[\frac{x -25.4}{13}=\frac{x-23.2}{24} = \frac{x-?}{17}\]

OpenStudy (anonymous):

The answer should be 24.6 cm. The rate of burning is 0.2 cm/hr, as the candle burns 2.2 cm in 11 hours. 2.2cm/11hrs = 0.2 cm/hr. After 4 hours (after the end of the thirteenth hour), the candle should have burned 0.8cm. Therefore the answer is 25.4cm - 0.8cm = 24.6cm.

OpenStudy (anonymous):

Find how much it is burning per hour. You know that in the 11 extra hours (24 minus 13), it burned 2.2cm (25.4 minus 23.2). So, per hour, it burns 2.2/11 = 0.2cm. So, in 4 extra hours (17 minus 13), it would burn 0.8cm. So, it would be of size 24.6cm (25.4 minus 0.8).